Alphagirl:

Can u help me with my 3 math pages

Alphagirl:

Alphagirl:

Alphagirl:

Sorry i ment 2 not 3

Anthony2005:

17 is not proportional

Alphagirl:

How is it not proportional

Anthony2005:

if it was proportional then it would be for example it would be up one over one something like that.

Alphagirl:

Ok

surjithayer:

17 is proportional constant of proportionality= (24-0)/(2-0)=24/2=12

Alphagirl:

Thanks

surjithayer:

20. a. y=22x b. for every additional acre the number of bushels increases by 22 i.e., if farm is of 1 acre ,number of bushels=22 for 2 acres,number of bushels=44 etc. c. size of farm number of bushels a b 0 0 10 220 25 550 (22*25=550) 30 660 (22*30=660) 50 1100 (22*50=1100) d. some coordinates are mentioned above. put x=0,10,20,30,...,90,100 y=0,220,440,660,...,1980,2200 draw the graph. e. (25,550) means when farm is of size 25 then number of bushels=550